Schueale000:

A different car can travel 30.5 miles for every gallon of gas used. Use proportional reasoning to find the value of x that completes the table showing this relationship.

dude:

Please make sure to include the table next time 

dude:

We're trying to find x here You want the same value type on the same side of the fraction You can have \(\dfrac{miles}{gallons}=\dfrac{miles}{gallons}\) In each fraction you want to have the corresponding value For x miles, we were given 3 gallons. So you can substitute that into one "fraction" \(\dfrac{x~mi}{3~gallons}=\dfrac{miles}{gallons}\) For the second "fraction" you can choose any point given in the table